OpenStudy (loser66):

Let \(f:\mathbb R^n\rightarrow \mathbb R^m\) and suppose there is a constant M such that for \(x\in \mathbb R^n,~~||f(x)||\leq M||x||^2\) Prove that f is differentiable at \(x_0=0\) and that \(Df(x_0)=0\) Please, help

OpenStudy (zzr0ck3r):

which norm is this?

OpenStudy (loser66):

vector space, I think!! That's all I have from the problem.

OpenStudy (zarkon):

I assume you mean \(x_o=(0,0,\cdots,0)\) (an n-tuple) clearly \(f(x_0)=0\) since \[\|f(x_0)\|\leq M\|x_0\|^2=0\] then just look at \[\frac{\|f(x)-f(x_0)\|}{\|x-x_0\|}=\frac{\|f(x)\|}{\|x\|}\leq M\|x\|\] and let \(x\to x_0\)

OpenStudy (loser66):

Thank you so much. I think I get it. :)